OpenStudy (anonymous):

Which polynomial is a perfect square trinomial? 36x2 − 18x − 9 4a2 − 28a − 49 25b2 − 40b + 16 4x2 − 6x + 9

OpenStudy (campbell_st):

a perfect square is \[(a + b)^2 = a^2 + 2ab + b^2\] a perfect square has the 1st and last terms positive... as squaring any negative gives a positive. and notice the middle term is double the product of a and b. hope this helps...

OpenStudy (anonymous):

eh

OpenStudy (campbell_st):

since the middle term in each is negative then \[(a -b)^2 = a^2 - 2ab + b^2\] is the expansion...

OpenStudy (campbell_st):

ok... for the 1st choice are the 1st and last terms positive..?

OpenStudy (anonymous):

no..

OpenStudy (campbell_st):

ok... so the same logic can be applied to the 2nd choice... on now the 3rd choice using the expansion above you are beaing with 5b and -4... is the middle term.. double the product of 5b and - 4 for the last choice you are looking at 2x and -3... is the middle term double the product of 2x and - 3..?

OpenStudy (anonymous):

so 25b2 − 40b + 16?

OpenStudy (campbell_st):

thats the correct solution... well done... I forgot to say \[25b^2 = (5b)^2 ...and....16 = (-4)^2\] the reason I know its -4 is that the middle term is negative.